when Christine was born her mother was 25 years old. Three years ago her mother was 6 times as old as Christine. Find Christine's present age.
Answer by ankor@dixie-net.com(22740)

Let c = Chis's age
let m = Mom's age
:
Write an equation for each statement, simplify as much as possible
:
when Christine was born her mother was 25 years old.
m = c + 25
:
Three years ago her mother was 6 times as old as Christine.
m - 3 = 6(c - 3)
m - 3 = 6c - 18
m = 6c - 18 + 3
m = 6c - 15
Replace m with with (c+25) from the 1st statement
c + 25 = 6c - 15
25 + 15 = 6c - c
40 = 5c
c = 40/5
c = 8 yrs is Chris's age
:
:
Check this:
Find Mom's age and check the solutions in the 2nd statement
"Three years ago her mother was 6 times as old as Christine."
m = 8 + 25 = 33 yrs
then
33 - 3 = 6(8 - 3)
30 = 6(5)